My Project

Generated by Doxygen 1.8.13

Contents

Index  1

Index

∼BMASocket BMASocket, 39

accept BMATCPServerSocket, 48

BMAAccount, 5 BMAAuthenticate, 6 BMACommand, 7 BMAConsoleServer, 8 BMAConsoleSession, 10

output, 11 BMAEPoll, 12

registerSocket, 14 unregisterSocket, 14

BMAFile, 15 BMAHTTPRequestHandler, 17 BMAHTTPServer, 18 BMAHTTPSession, 20

onDataReceived, 21 BMAHeader, 16 BMALog, 22 BMAMP3File, 22 BMAMP3StreamContentProvider, 23 BMAMP3StreamFrame, 24 BMAObject, 26 BMAParseHeader, 26 BMAProperty< T >, 26 BMASIPINVITE, 29 BMASIPREGISTER, 30 BMASIPRequestHandler, 31 BMASIPServer, 33 BMASIPSession, 35

onDataReceived, 36 BMASession, 27

onConnected, 28 onDataReceived, 28 output, 29

BMASocket, 37 ∼BMASocket, 39 eventReceived, 39 onConnected, 39 onDataReceived, 39 onRegistered, 40

BMAStreamContentProvider, 40 BMAStreamFrame, 41 BMAStreamServer, 42 BMAStreamSession, 44

onDataReceived, 45 BMATCPServerSocket, 46 accept, 48 onDataReceived, 48

BMATCPSocket, 48 onConnected, 50 output, 50 BMAThread, 50 BMATimer, 51

onDataReceived, 52 BMAUDPServerSocket, 53 onDataReceived, 54

BMAUDPSocket, 55

eventReceived BMASocket, 39

onConnected BMASession, 28 BMASocket, 39 BMATCPSocket, 50

onDataReceived BMAHTTPSession, 21 BMASIPSession, 36 BMASession, 28 BMASocket, 39 BMAStreamSession, 45 BMATCPServerSocket, 48 BMATimer, 52 BMAUDPServerSocket, 54

onRegistered

BMASocket, 40 output

BMAConsoleSession, 11 BMASession, 29 BMATCPSocket, 50

registerSocket BMAEPoll, 14

unregisterSocket BMAEPoll, 14